How to solve -5x+12=-x-9
Answer by ad_alta(240)

1) -5x+12=-x-9 (add 5x to both sides)
2) 12=4x-9 (add 9 to both sides)
3) 21=4x (divide both sides by 4)
4) x=21/4
5) celebrate!